# Beyond Diagnosis: Evaluating Multimodal LLMs for Pathology Localization in Chest Radiographs

### v1

@@ 5 Discussion
Beyond diagnosis, we evaluated the ability of MLLMs to localize findings in chest radiographs. We found that performance was mixed, with GPT-4 and GPT-5 outperforming a domain-specific CNN baseline on two pathologies, but underperforming the CNN and human benchmark overall. Notably, these two pathologies tend to appear in consistent anatomical locations. Nonetheless, GPT-5 showed strong improvements compared to GPT-4 on the more spatially variable pathologies, and its predictions were largely anatomically plausible even when it did not precisely localize the pathology. MedGemma, a medicine-specific model that includes chest radiographs in its training, exhibited the lowest performance overall and had a high proportion of anatomically-implausible predictions.
Our findings have several implications for clinical practice and AI development. First, while (M)LLMs have recently shown strong performance on question-answer challenges and diagnostic tasks (Eriksen et al., 2023; Suh et al., 2024; Nori et al., 2023), our results show that leading models struggle with fine-grained spatial reasoning. As a smaller model, MedGemma may especially be limited in the ability to generalize to novel tasks compared to traditional benchmarks used in the field. However, the improvements in accuracy and medical understanding observed with GPT-5 compared to GPT-4 are encouraging. While it remains to be seen whether these gains will continue to occur with future model versions, these results at least suggest that large, general-purpose MLLMs are a promising approach compared to domain-specific vision-language models. In the meantime, agentic strategies that integrate the flexibility of LLMs with task-specific tools may be the best strategy, where recent work in the context of chest radiograph interpretation has demonstrated this possibility (Fallahpour et al., 2025).
@@
Our study has several limitations. Our analysis focused on one dataset given the need for localized annotations and meaningful benchmark comparisons. Future work could extend our analysis pipeline to other datasets and tasks. Furthermore, some pathologies in the CheXlocalize dataset have a small image count, leading to large error bars for the results for those pathologies. Another challenge lies in the variability of ground truth annotation sizes across images and pathologies, which makes it difficult to define a hit rate metric that robustly captures all scenarios. Additionally, while interpretable, hit rate metrics sacrifice the flexibility and granularity offered by other localization metrics, such as IoU. Finally, although our goal was to assess default out-of-the-box performance on the proposed task, more extensive prompt engineering could be explored in the future.
@@ Conclusion
Prompt

### v2

@@ 5 Discussion
Beyond diagnosis, we evaluated the ability of frontier MLLMs to localize findings in chest radiographs. We found that performance was mixed, with GPT-4 and GPT-5 performing relatively strong on pathologies that appear in consistent anatomical locations, but underperforming the CNN and human benchmark overall. Nonetheless, GPT-5 showed strong improvements compared to GPT-4 on the more spatially variable pathologies, and its predictions were largely anatomically plausible even when it did not precisely localize the pathology. MedGemma, a medicine-specific model that includes chest radiographs in its training, exhibited the lowest performance overall and had a high proportion of anatomically-implausible predictions. Few shot prompting brought MedGemma’s performance closer to the other models, with modest improvements for GPT-4 and GPT-5 as well.
Our findings have several actionable implications for clinical practice and AI development. First, while (M)LLMs have recently shown strong performance on question-answer challenges and diagnostic tasks (Eriksen2023-ye; Suh2024-qw; nori2023capabilitiesgpt4medicalchallenge), our results show that leading models struggle with fine-grained spatial reasoning. Smaller models such as MedGemma (and Gemma 3 27B on which it was built) may especially be limited in the ability to generalize to novel tasks, such as the one introduced here, compared to traditional benchmarks used in the field. Thus, the observed results may be partly attributable to the task itself rather than inherent medical knowledge, highlighting the importance of rigorous evaluation and adaptation strategies (e.g., few shot prompting) for specific use cases in clinical practice. While GPT-5’s performance remained well below the human benchmark, the improvements in accuracy and medical understanding observed with GPT-5 compared to GPT-4 are encouraging. It remains to be seen whether these gains will continue to occur with future model versions, but these results at least suggest that large, general-purpose MLLMs are a promising approach compared to domain-specific vision-language models. In the meantime, agentic strategies that integrate the flexibility of LLMs with task-specific tools may be the best strategy, where recent work in the context of chest radiograph interpretation has demonstrated this possibility (medrax).
@@
Our study has several limitations. Our analysis focused on one dataset given the need for localized annotations and meaningful benchmark comparisons. Future work could extend our analysis pipeline to other datasets and tasks. Furthermore, some pathologies in the CheXlocalize dataset have a small image count, leading to large error bars for the results for those pathologies. We included these pathologies because they still reveal meaningful model behavior, as illustrated in Fig. 5 where GPT-4 and GPT-5 localize a lung pathology to the shoulder. Another challenge lies in the variability of ground truth annotation sizes across images and pathologies, which makes it difficult to define a hit rate metric that robustly captures all scenarios. Additionally, while interpretable, hit rate metrics sacrifice the flexibility and granularity offered by other localization metrics, such as IoU. Finally, although we explored different prompting strategies, more extensive prompt engineering could be explored in the future.
@@ Conclusion
[…]
W.L. acknowledges funding support from NIBIB award R21EB035247 and NLM award R01LM014775.
@@
Prompts
Baseline zero shot prompt
@@
Chain-of-Thought Prompt
System: You are an expert chest radiologist specializing in analyzing view chest X-rays. Your task is to precisely localize abnormalities using a grid overlay. Approach this task methodically, using your expertise to analyze the image step-by-step.
User: [Image]
Your task:
  1. 1.
    Carefully examine the entire chest X-ray image, paying close attention to areas where ‘{condition}’ typically manifests.
  2. 2.
    Identify all grid cells where you can observe signs of ‘{condition}’. List these cells in your thought process.
  3. 3.
    For each identified cell, briefly note the specific features or abnormalities you observe that are consistent with ‘{condition}’.
  4. 4.
    Compare the identified cells and their features. Consider which cell contains the most prominent or representative manifestation of ‘{condition}’.
  5. 5.
    If the condition spans multiple cells, determine which single cell encompasses the most significant part of the abnormality.
  6. 6.
    Based on your analysis, select the single grid coordinate (e.g., D5) that best represents the location of ‘{condition}’ in this X-ray.
  7. 7.
    Provide your final answer as a single grid coordinate without any additional explanation or text.
Example thought process: “I observe signs of ‘{condition}’ in cells C4, D4, and D5. In C4, there’s [specific feature]. D4 shows [another feature]. D5 contains [feature]. Comparing these, D5 appears to be the most representative location for ‘{condition}’. Final answer: D5”
Fewshot Prompt
System: You are an expert chest radiologist specializing in analyzing chest X-rays. You are given a gridded {view} image of a chest X-ray with the abnormality ‘{condition}’ confirmed to be present.
User: Zeroshot Prompt with [Example Image 1]
Assistant: Ground Truth grid coordinate for Example 1
User: Zeroshot Prompt with [Example Image 2]
Assistant: Ground Truth grid coordinate for Example 2
User: Zeroshot Prompt with [Example Image 3]
Assistant: Ground Truth grid coordinate for Example 3
User: Zeroshot Prompt with [Test Image]
Few shot selection: Three random examples from the validation split with the same view position and pathology were selected. If there were not enough examples in the validation split, the remainder were selected randomly from the test split (excluding the image being evaluated).
@@
Table 2: Hit rate per pathology for each prompting strategy. Bold indicates the highest performance for the pathology; underline indicates the second highest.
Model Prompt
Airspace
Opacity
Atelect. Cardiomeg. Consol. Edema E. Cardio.
Lung
Lesion
Pleural
Effusion
Pneumo.
Random Baseline - 9.8 9.5 16.0 10.4 18.1 20.2 6.8 8.6 7.5
[…]
MedGemma Zero Shot 12.3 9.0 42.3 14.3 28.9 28.6 14.3 10.0 0.0
CoT 6.1 7.9 37.7 17.1 31.3 83.2 0.0 9.2 10.0
Few Shot 23.3 23.7 57.7 37.1 43.4 62.6 7.1 23.3 10.0
[…]
CoT 36.2 44.6 72.6 31.4 61.4 91.9 21.4 36.7 30.0
[…]
Gemma3 27B Zero Shot 12.0 11.3 19.4 17.1 20.5 53.9 7.1 7.5 10.0
CoT 6.1 9.6 33.1 17.1 50.6 83.5 21.4 3.3 0.0
Few Shot 16.8 21.5 42.9 5.7 34.9 73.1 14.3 20.0 0.0
Image Count - 309 177 175 35 83 297 14 120 10